[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Jul 08, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Fri Jul 8 13:56:50 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:

HSU:
1. ATCF: The Pacific desk ATCF workstation (pac-atcf.nhc.noaa.gov) remains  
sluggish and the Firefox web browser on that machine has no access to the  
internet. NCO Linux SysAdmins will downgrade the Linux kernel to remedy  
these problems.

2. ATCF: No fixes from TAFB were imported at 22:46 UTC yesterday. ATCF  
dataflow expert Monica Bozeman will investigate.
Other items:
1. Another kernel panic occurred on the NESDIS/STAR primary ftp server  
(ftp.star.nesdis.noaa.gov) yesterday due to a memory leak in the  
THREDDS/TomCat server, which caused outages in the Himawari-8 satellite  
data. Users were advised to use the backup server  
(ftp2.star.nesdis.noaa.gov) for data retrieval. Service was restored about  
4 hours later.

2. A GOES-13 (GOES-EAST) satellite maneuver is scheduled for 15:33-15:44  
UTC on Tuesday, July 19. NESDIS claims that there will be no imager or  
sounder impacts because the maneuver operations will be performed during  
the daily Housekeeping period.

3. Earth Networks Incorporated (ENI) lightning data products stopped  
updating in AWIPS yesterday morning at 13:45 UTC. NCF was not receiving  
lightning data from Line 205. Earth Networks tech support was contacted by  
the NOAA-NWS Telecommunication Operations Center (TOC) to restore the data  
flow.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160708/ec83786b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list